OpenStudy (anonymous):

Which pairs are solutions to the equation? 5xy + 9 = 44 (Points : 1) (1, 7) and (4, 11) (1, 7) and (7, 1) (7, 1) and (2, 3) (4, 11) and (2, 3)

OpenStudy (anonymous):

xy=7 kind of weird.

OpenStudy (anonymous):

y=7/x

OpenStudy (anonymous):

so I choose b

OpenStudy (haseeb96):

5xy+9=44 5xy=44-9 5xy=35 xy=35\5 xy=7 ,answer is b. let see how (1,7) in it x=1 and y=7 5(1)(7)+9=44 35+9=44 44=44 hence one orderpair is satisfied take another (7,1) x=7 and y=1 put it in the given equation 5(7)(1)+9=44 35+9=44 44=44 hence shown

OpenStudy (anonymous):

thank you